OpenStudy (anonymous):

Prove \[\lim_{x \rightarrow 0} \sqrt{4-x} = 2\]using \(\epsilon\) - \(\delta\) definition.

OpenStudy (anonymous):

For δ one: 0 < x-0 < δ => 0< x <δ |f(x) - L| < ϵ | \(\sqrt{4-x}\) - 2 | < ϵ - ϵ < \(\sqrt{4-x}\) - 2 < ϵ 2-ϵ < \(\sqrt{4-x}\) < ϵ + 2 (2-ϵ)^2 < 4-x < (ϵ + 2)^2 ϵ^2 - 4ϵ < -x < ϵ^2 + 4ϵ 4ϵ - ϵ^2 > x > -ϵ^2 - 4ϵ Then, I'm lost!

zepdrix (zepdrix):

 Ummmmm I don't think you want to square each term like that, I think it'll get you into more trouble D: I think you wanna go about it like this somehow... lemme do a few more steps.. see if i remember how to do this correctly.